OpenStudy (anonymous):

Help with 5 questions please. Will give fan and medal

OpenStudy (anonymous):

Select the coordinates of two points on the line y = -2. (2, -2) and (-2, 2) (-2, -2) and (-2, 0) (2, -2) and (0, -2) (-2, 2) and (-2, -2)

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7

OpenStudy (anonymous):

Is the line for the equation y = 1 horizontal or vertical? What is the slope of this line? Select the best answer. The line is horizontal. The slope is 0. The line is horizontal. The slope is 1. The line is vertical. The slope is undefined. The line is vertical. The slope is 0.

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

C A C D

OpenStudy (anonymous):

3rd question's answer is : Horizontal, gradient 0

OpenStudy (anonymous):

wow thank you

OpenStudy (anonymous):

how about the 5th one though

OpenStudy (anonymous):

@ArsalanAijaz

OpenStudy (anonymous):

I have answered 1,2,4,5 in a row. Then 3rd

OpenStudy (anonymous):

Sorry, I just skipped 3rd by mistake earlier

Nnesha (nnesha):

@ArsalanAijaz welcome to open study!!!! http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Am I violating it?

OpenStudy (anonymous):

I will explain my answers too, I'm not running away :D

Nnesha (nnesha):

@hi10500 post one question at a time learn the method then try do rest of them by ur own i know you can do it so try it :)

OpenStudy (anonymous):

are you sire about all of these @ArsalanAijaz

OpenStudy (anonymous):

*sure

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I skipped the third so I answered it later. If you can stay, I'll explain

OpenStudy (anonymous):

who do you know all of them you literally told me all of them in like 2 minutes

OpenStudy (anonymous):

I just do maths :D

OpenStudy (anonymous):

Plus, with practice, you just get faster and faster

OpenStudy (anonymous):

ok well the third one is horizontal slope 0 right

OpenStudy (anonymous):

@ArsalanAijaz

Directrix (directrix):

@ArsalanAijaz OpenStudy values the Learning process - not the ‘Give you an answer’ process' Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Yup. Cuz Y= something means that it is horizontal cuz y1-y2 would always give 0 and 0 divided by anything is 0

OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1 YOU

OpenStudy (anonymous):

@ArsalanAijaz

OpenStudy (anonymous):